OpenStudy (anonymous):

Select the ordered pair from the choices below that is a solution to the following system of equations:5y = x - 7 4x - y = -10

OpenStudy (unklerhaukus):

substitute the co-ordinates into each potential solution into the system if the resulting equations are true then that particular orded pair solves the system

OpenStudy (anonymous):

i dont think i did it right i meant Select the ordered pair from the choices below that is a solution to the following system of equations:5y = x - 7 4x - y = -10

OpenStudy (anonymous):

help please

OpenStudy (unklerhaukus):

./?where are the choices

OpenStudy (anonymous):

a. (-8,3) b. (1/5,8) c. (-3,-2) d.(-1,-1)

OpenStudy (anonymous):

i got c idk if im right ? correct me if im wrong

OpenStudy (unklerhaukus):

are both these equations true? 5(-2) = (-3) - 7 4(-3) - (-2) = -10